# BiLLM: Pushing the Limit of Post-Training Quantization for LLMs

### v1

@@ Abstract
Pretrained large language models (LLMs) exhibit exceptional general language processing capabilities but come with significant demands on memory and computational resources. As a powerful compression technology, binarization can extremely reduce model weights to a mere 1 bit, lowering the expensive computation and memory requirements. However, existing quantization techniques fall short of maintaining LLM performance under ultra-low bit-widths. In response to this challenge, we present BiLLM, a groundbreaking 1-bit post-training quantization scheme tailored for pretrained LLMs. Based on the weight distribution of LLMs, BiLLM first identifies and structurally selects salient weights, and minimizes the compression loss through an effective binary residual approximation strategy. Moreover, considering the bell-shaped distribution of the non-salient weights, we propose an optimal splitting search to group and binarize them accurately. BiLLM achieving for the first time high-accuracy inference (e.g. 8.41 perplexity on LLaMA2-70B) with only 1.08-bit weights across various LLMs families and evaluation metrics, outperforms SOTA quantization methods of LLM by significant margins. Moreover, BiLLM enables the binarization process of the LLM with 7 billion weights within 0.5 hours on a single GPU, demonstrating satisfactory time efficiency. The code is available at https://github.com/Aaronhuang-778/BiLLM.
Model Binarization, Large Language Model, Model Compression, Deep Learning
@@ 1 Introduction
Recently, large language models (LLMs) based on transformers (Vaswani et al., 2017) have garnered significant attention in natural language processing. Pretrained LLMs like OPT (Zhang et al., 2022) and LLaMA (Touvron et al., 2023a), have demonstrated excellent performance across a range of evaluation benchmarks. However, LLMs pose substantial challenges in deployment on memory-constrained devices due to their immense parameter size and computation requirements. For instance, the widely-used LLaMA2-70B (Touvron et al., 2023b) model, with its 70 billion parameters, requires 150 GB of storage in half-precision (FP16) format. This necessitates at least two A100 GPUs, each with 80 GB of storage space, for inference.
Figure 1: The perplexity of LLaMA-13B on WikiText2 under different bit-widths. Round-to-nearest (RTN), GPTQ, and PB-LLM (10% weight of INT8) suffer accuracy loss at ultra-low bits, facing the sharply increasing perplexity ($\downarrow$). BiLLM demonstrates exceptional performance under binarization.
@@ 2.2 Network Binarization
$\vspace{-0.1in}\mathbf{W}_{b}=\alpha\cdot\operatorname{sign}(\mathbf{W}_{f}),$ (1)
@@ 3.1 Salient Weight Binarization for LLMs
where $\mathbf{B}\in\{-1,+1\}^{k\times m}$, $k$ is the number of selected columns. The problem (Rastegari et al., 2016) of optimal $\alpha$ and $\mathbf{B}$ can simply be solved as $\alpha=\frac{||\mathbf{W}||_{\ell 1}}{m}$ and $\mathbf{B}=\operatorname{sign}(\mathbf{W})$. Then, the optimization function for selecting salient columns is defined as:
@@ 3.3 Pipeline of BiLLM
Figure 6: Weights and hardware overhead changes on Llama-7B. The left picture shows the calculation parameters as a function of the significant weight ratio; the right picture shows the hardware overhead as a function of the block.
Table 1: Average bit results from structural searching and residual binarization of OPT, LLaMA, and LLaMA2 families.
@@ 4.2 Results
Figure 7: GPTQ, PB-LLM, BiLLM performed on the PTB and c4 datasets, mainly on LLaMA-7B, LLaMA2-7B, and OPT-6.7B, and we found that BiLLM performed relatively well.
@@ 5 Conclusions
6 Impact Statements
@@ Appendix C Searching Curve of Salient Column and Non-salient Distribution
Figure 10 shows the unstructured search curve for the non-salient weights in the OPT6.7B model, with the same composition as that in Figure 9. The horizontal axis represents the ratio between $p$ and the maximum weight value. Despite searching on a block-wise basis, the search curve still exhibits convex properties, indicating the presence of an optimal $*p*$. This phenomenon demonstrates that the non-salient weights exhibit characteristics closely resembling an ideal Gaussian or Laplacian distribution (You, 2010; Fang et al., 2020).

### v2

@@ Abstract
Pretrained large language models (LLMs) exhibit exceptional general language processing capabilities but come with significant demands on memory and computational resources. As a powerful compression technology, binarization can extremely reduce model weights to a mere 1 bit, lowering the expensive computation and memory requirements. However, existing quantization techniques fall short of maintaining LLM performance under ultra-low bit-widths. In response to this challenge, we present BiLLM, a groundbreaking 1-bit post-training quantization scheme tailored for pretrained LLMs. Based on the weight distribution of LLMs, BiLLM first identifies and structurally selects salient weights, and minimizes the compression loss through an effective binary residual approximation strategy. Moreover, considering the bell-shaped distribution of the non-salient weights, we propose an optimal splitting search to group and binarize them accurately. BiLLM, for the first time, achieves high-accuracy inference (e.g. 8.41 perplexity on LLaMA2-70B) with only 1.08-bit weights across various LLM families and evaluation metrics, outperforms SOTA quantization methods of LLM by significant margins. Moreover, BiLLM enables the binarization process of a 7-billion LLM within 0.5 hours on a single GPU, demonstrating satisfactory time efficiency. Our code is available at https://github.com/Aaronhuang-778/BiLLM.
Machine Learning, ICML
@@ 1 Introduction
Recently, large language models (LLMs) based on transformers (Vaswani et al., 2017) have garnered significant attention in natural language processing. Pre-trained LLMs such as OPT (Zhang et al., 2022) and LLaMA (Touvron et al., 2023a), have demonstrated excellent performance across a range of evaluation benchmarks. However, LLMs pose substantial challenges in deployment on memory-constrained devices due to their immense parameter size and computation requirements. For instance, the widely-used LLaMA2-70B (Touvron et al., 2023b) model, with its 70 billion parameters, requires 150 GB of storage in half-precision (FP16) format. This necessitates at least two A100 GPUs, each with 80 GB of storage space, for inference.
Figure 1: Perplexity of LLaMA-13B on WikiText2 under different bit-widths. Round-to-nearest (RTN), GPTQ, and PB-LLM (10% weight of INT8) suffer accuracy loss at ultra-low bits, facing the sharply increasing perplexity ($\downarrow$). BiLLM demonstrates exceptional performance under binarization.
@@ 2.2 Network Binarization
$\mathbf{W}_{b}=\alpha\cdot\operatorname{sign}(\mathbf{W}_{f}),$ (1)
@@ 3.1 Salient Weight Binarization for LLMs
where $\mathbf{B}\in\{-1,+1\}^{n\times k}$, $k$ is the number of selected columns. The problem (Rastegari et al., 2016) of optimal $\alpha$ and $\mathbf{B}$ can simply be solved as $\alpha=\frac{||\mathbf{W}||_{\ell 1}}{n\times k}$ and $\mathbf{B}=\operatorname{sign}(\mathbf{W})$. Then, the optimization function for selecting salient columns is defined as:
@@ 3.3 Pipeline of BiLLM
Figure 6: Weights and hardware overhead changes on Llama-7B. The left shows the calculation parameters as a function of the significant weight ratio; the right shows the hardware overhead as a function of the block.
Table 1: Average bit results from structural searching and residual binarization of OPT, LLaMA, and LLaMA2.
@@ 4.2 Results
Figure 7: GTPQ, PB-LLM, BiLLM performed on the PTB and c4 datasets, mainly on LLaMA-7B, LLaMA2-7B, and OPT-6.7B, and we found that BiLLM performed relatively well.
@@
Table 5: Model size comparison of LLaMA family.
Method LLaMA-7B LLaMA2-7B LLaMA-13B LLaMA2-13B LLaMA-30B LLaMA-65B LLaMA-70B
FP16 13.5GB 13.5 GB 24.2 GB 25.0 GB 60.5 GB 121.0 GB 129.3 GB
BiLLM 1.5 GB 1.6 GB 2.7 GB 2.8 GB 6.1 GB 14.8 GB 15.4 GB
@@
Table 6: The memory occupancy rate compared with FP16 and the corresponding accuracy on OPT-30B.
Configuration BiLLM PB-LLM (10%) GPTQ
Average bit-width 1.11 1.7 2
Memory Occupancy 9.70% 16.50% 13.30%
PPL on WikiText-2 12.71 25.14 15.71
  • *: Equation of memory occupancy [R6]: memory_occupancy_compared_w_FP16 = (binary_unsalint_weight_size + residual_binary_salint_weight_size + CSR_compressed_bitmap_size + scaling_factor_size) / floating_point_weight_size.
Table 7: Memory occupancy rate compared with FP16 OPT.
      Model       BiLLM       GPTQ-2bit
      OPT-1.3B       9.40%       13.30%
      OPT-2.7B       9.50%       13.30%
      OPT-6.7B       9.30%       13.30%
      OPT-13B       9.70%       13.30%
      OPT-30B       9.70%       13.30%
Model size. In Table 5, we present the FP16 of models ranging from LLaMA-7B to 65B and LLaMA2-7B to 70B, as well as the sizes of binarized models after compression by BiLLM. Notably, BiLLM achieved close to a tenfold compression of weights across LLMs of different sizes.
GPU memory. The motivation of our BiLLM is to push the bit-width compression limit of LLM weights under post-training conditions, which reduces both the storage and GPU memory footprint of LLMs and retains their accuracy to the greatest extent for being practical. Although binarized GEMM is hard to implement directly due to fine-grained grouping, the extreme bit-width compression of our BiLLM brings significant savings in GPU memory requirements (size and bandwidth), which is considered to be one of the most significant efficiency bottlenecks of LLM inference (Gholami et al., 2024; Dettmers et al., ; Dettmers et al., 2023a; Xiao et al., 2023; Chee et al., 2024; Shang et al., 2023). Here, we provide detailed memory and performance comparisons to demonstrate the advantages of BiLLM (as shown in Table A.1): for the OPT-30B model, BiLLM (1.1-bit) achieves a 41.57% and 27.07% memory compression improvement compared to PB-LLM (1.7-bit) and GPTQ (2-bit), respectively, while enhancing accuracy by 49.44% and 19.10%. We further provide a detailed comparison of memory usage with the 2-bit GPTQ method under different sizes of LLM in Table 7. The memory occupancy of our BiLLM is only about 69.9% of 2-bit quantization, which shows the great memory-saving benefit of our BiLLM from the extreme bit-width reduction, and we also achieve higher accuracy with the significantly saved memory.
@@ 5 Conclusions
Acknowledgement. This work was supported by the National Science and Technology Major Project (2021ZD0110503), the Swiss National Science Foundation (SNSF) project 200021E_219943 Neuromorphic Attention Models for Event Data (NAMED), the Baidu Scholarship, and the National Natural Science Foundation of China (No. 62306025, No. 92367204).
Impact Statement
@@ Appendix C Searching Curve of Salient Column and Non-salient Distribution
Figure 10 shows the unstructured search curve for the non-salient weights in the OPT6.7B model, with the same composition as that in Figure 9. The horizontal axis represents the ratio between $p$ and the maximum weight value. Despite searching on a block-wise basis, the search curve still exhibits convex properties, indicating the presence of an optimal $p*$. This phenomenon demonstrates that the non-salient weights exhibit characteristics closely resembling an ideal Gaussian or Laplacian distribution (You, 2010; Fang et al., 2020).